# Look Before You Fuse: 2D-Guided Cross-Modal Alignment for Robust 3D Detection

### v1

@@ Abstract
Integrating LiDAR and camera inputs into a unified Bird’s-Eye-View (BEV) representation is crucial for enhancing 3D perception capabilities of autonomous vehicles. However, current methods are often affected by misalignment between camera and LiDAR features. This misalignment leads to inaccurate depth supervision in camera branch and erroneous fusion during cross-modal feature aggregation. The root cause of this misalignment lies in projection errors, stemming from minor extrinsic calibration inaccuracies and rolling shutter effect of LiDAR during vehicle motion. In this work, our key insight is that these projection errors are predominantly concentrated at object-background boundaries, which are readily identified by 2D detectors. Based on this, our main motivation is to utilize 2D object priors to pre-align cross-modal features before fusion. To address local misalignment, we propose Prior Guided Depth Calibration (PGDC), which leverages 2D priors to correct local misalignment and preserve correct cross-modal feature pairs. To resolve global misalignment, we introduce Discontinuity Aware Geometric Fusion (DAGF) to process calibrated results from PGDC, suppressing noise and explicitly enhancing sharp transitions at object-background boundaries. To effectively utilize these transition-aware depth representations, we incorporate Structural Guidance Depth Modulator (SGDM), using a gated attention mechanism to efficiently fuse aligned depth and image features. Our proposed method achieves state-of-the-art performance on nuScenes validation dataset, with its mAP and NDS reaching 71.5% and 73.6% respectively.
Figure 1: Projection errors misplace distant points from the wall (cyan/green) onto the foreground car (which should be red). In contrast, background points, like those on the wall (orange) and garage (blue), are projected correctly, clearly defining the boundary between them. This highlights the system’s failure to handle sharp foreground-background depth changes, which can be efficiently captured by 2D detectors.
@@ Introduction
Robust 3D perception is a cornerstone of autonomous driving, where the effective fusion of complementary sensors like cameras and LiDAR is essential. For instance, images provide rich semantic representations but lack accurate depth information. Conversely, point clouds deliver precise geometric and depth cues but suffer from sparsity and absence of semantic context. Thus, effectively harnessing the complementary strengths of multi-modal data while addressing their inherent limitations is critical for building robust perception systems.
While promising, the fusion of LiDAR and camera data introduces significant challenges, primarily stemming from inherent cross-modal misalignment. Specifically, the inherent misalignment between point cloud and image creates obstacles that undermine the effectiveness of multi-sensor fusion (Yu et al. 2023). For example, a fundamental challenge in BEV perception lies in recovering the implicit depth information from 2D images to generate precise BEV feature representations for downstream neural networks. And therefore, a variety of approaches have focused on leveraging LiDAR signals to provide explicit depth supervision (Li et al. 2023) or develop implicit fusion methods that integrate camera and LiDAR information to implicitly provide depth information (Liu et al. 2023; Liang et al. 2022). Nevertheless, the misalignment of point cloud and image features due to calibration errors remains an insufficiently addressed problem, which leads to two major issues. First, the misalignment between the camera and LiDAR sensors leads to missing or erroneous depth supervision, degrading depth estimation accuracy, especially in areas of dramatic depth variation. Second, during the information fusion process, incorrect geometry-camera pairings can confuse the fusion module, making effective integration particularly challenging.
@@
Existing approaches have attempted to mitigate cross-modal misalignment, yet each suffers from significant drawbacks. To ensure feature consistency, some methods like ObjectFusion (Cai et al. 2023) adopt an object-centric strategy to avoid direct projection errors. While this effectively sidesteps the main cause of misalignment, it comes at the cost of sacrificing crucial contextual information. Other approaches like MetaBEV (Ge et al. 2023) and RobBEV (Wang et al. 2024) attempt to mitigate misalignment effects by designing more robust and adaptive fusion modules. Although these modules are more resilient to inconsistent inputs, they cannot correct the initial geometric errors from the view transformation. In essence, they are skillfully fusing features that have already been misplaced. Finally, global alignment techniques like GraphBEV (Song et al. 2024a) directly address the geometric problem, effectively eliminating misalignment in areas with steep depth gradients. However, they tend to unnecessarily smooth geometrically stable regions where misalignment is negligible or even entirely absent, thereby incorrectly modifying already-correct depth values.
These methods ignore the fact that misalignment stems from two primary sources, both of which cause projection errors that are amplified with increasing depth: minor extrinsic calibration errors and motion-induced distortions from the LiDAR’s rolling shutter effect. These projection errors are minor for points on nearby objects but become significant for points on distant background, creating a severe misalignment of features at their shared boundaries. Targeting the root causes of misalignment, we propose a framework that uses 2D priors to guide geometric correction and mitigate misalignment. Our core insight is that critical errors predominantly occur at the boundaries between foreground objects and background regions where depth changes sharply, as illustrated in Figure 1.
@@
In this paper, we introduce three synergistic modules. First, to address the misalignment problem at its root, we propose (1) Prior Guided Depth Calibration (PGDC). Our core philosophy, “Look Before You Fuse”, proposes that instead of attempting to sort out already-misaligned data, we should in advance correct it using high-level semantic guidance. PGDC actively uses 2D detection proposals as explicit geometric priors to locate and correct misaligned point cloud data, providing a significantly more accurate depth map to supervise the camera branch’s depth estimation. However, a corrected sparse map alone is not enough. To enable the network to fully understand the scene’s geometric structure, we introduce (2) Discontinuity Aware Geometric Fusion (DAGF). Rather than using the raw depth map, DAGF processes it to create a dense and structurally aware representation by suppressing noise and explicitly enhancing sharp transitions at object boundaries. This enhanced representation, which encodes both reliable geometry and critical boundary information, is then fused with camera features, providing direct and unambiguous geometric guidance to reduce artifacts in final BEV output. To effectively utilize these calibrated signals for the view transformation, we introduce the (3) Structural Guidance Depth Modulator (SGDM). This module intelligently fuses image features and dense geometric representation from DAGF using a gated attention mechanism. Its purpose is to predict a highly accurate depth distribution for each pixel, enabling a more precise projection of features into the final BEV space.
@@
Figure 2: Overview of our proposed framework. It introduces two core modules, Prior Guided Depth Calibration (PGDC) and Discontinuity Aware Geometric Fusion (DAGF), to proactively mitigate multi-sensor feature misalignment before view transformation. Fusing the calibrated camera BEV features with LiDAR BEV features leads to robust 3D detection.
@@
Figure 3: The Prior Guided Depth Calibration (PGDC) module leverages 2D detection boxes as priors to precisely target and correct the most severe feature misalignments, which are caused by calibration errors and motion distortion. By applying localized smoothing to the point cloud within these detected regions, the module corrects the erroneous depth supervision signal. Simultaneously, it enhances the features of these critical image areas.
Our contributions are summarized below:
@@
  • We reveal that feature misalignment predominantly occurs at object-background boundaries and propose Prior Guided Depth Calibration (PGDC), a novel module that actively uses 2D priors to locate and resolve such misalignment, thereby providing more accurate depth supervision. To our best knowledge, our work is the first to use 2D detection results as explicit guidance for correcting cross-modal geometric inconsistencies before fusion.
@@
  • We introduce the Structural Guidance Depth Modulator (SGDM), an efficient attention-based module that fuses calibrated visual and geometric cues to generate a highly accurate depth distribution for view transformation.
@@
As shown in Figure 2, our proposed framework is a multi-sensor fusion pipeline built upon the strong BEVFusion (Liang et al. 2022) baseline, designed to explicitly mitigate feature misalignment in Bird’s Eye View (BEV) perception. The framework processes inputs from a set of $N$ surround-view cameras and a 360-degree LiDAR sensor. Our contribution is a framework of three synergistic modules. First, Prior Guided Depth Calibration (PGDC) uses 2D bounding boxes to correct misaligned LiDAR points at object boundaries, outputting a refined sparse depth map and enhanced image features. Second, Discontinuity Aware Geometric Fusion (DAGF) uses the corrected depth to generate a dense representation that captures reliable geometric structure. Finally, the Structural Guidance Depth Modulator (SGDM) fuses the enhanced image features (from PGDC) and the dense geometric representation (from DAGF) to predict a depth distribution for each camera view. Following the Lift-Splat-Shoot (LSS) paradigm, these are projected into a unified BEV feature map, which is then fused with LiDAR BEV features to produce robust 3D detection results.
@@
As shown in Figure 3, the Prior Guided Depth Calibration (PGDC) module operates independently on each of the $N$ camera views to correct the initial sparse depth supervision derived from LiDAR. For each view $i\in\{1,...,N\}$, the inputs are the image features $F_{\text{img}}^{(i)}\in\mathbb{R}^{H\times W\times C}$ and the corresponding sparse depth map $D_{\text{raw}}^{(i)}\in\mathbb{R}^{H\times W}$, which is generated by projecting the LiDAR point cloud into that camera’s image plane.
First, a 2D detection head provides a set of bounding boxes $\{B_{j}^{(i)}\}$ for each image. Based on our observation that misalignment is concentrated at object boundaries, we use these boxes to isolate critical regions. For each box $B_{j}^{(i)}$, we filter the LiDAR points whose projections fall within it. To mitigate misalignment, we introduce a novel smoothing operation detailed in Algorithm 1. Instead of simple averaging, our method captures the local depth structure more effectively. For each LiDAR point projected to a pixel $p$ with depth $d_{p}$, we first use a KD-Tree to find its 10 nearest neighbors, $\mathcal{N}_{p}$. We then select the two neighbors with the smallest depth and the two with the largest depth, forming a set of four critical neighbors, $\mathcal{N}_{\text{critical}}\subset\mathcal{N}_{p}$. This selection strategy is designed to simultaneously capture the object’s own depth consistency (via the nearest points) and the sharp depth discontinuity at the boundary between the object and the background (via the farthest points), thereby preserving critical information while smoothing noise. The depth of the original point is concatenated with the depths of these four selected neighbors, creating a 5-channel feature map, $f_{p}$, for that pixel:
@@
This feature map is then processed through a lightweight convolutional block—consisting of a Convolution, Batch Normalization, and a ReLU activation—to produce the final single-channel, smoothed depth value:
@@
9:     $\mathcal{D}_{\text{critical}}\leftarrow\{\mathcal{D}_{\text{sorted}}[0],\mathcal{D}_{\text{sorted}}[1],\mathcal{D}_{\text{sorted}}[-2],\mathcal{D}_{\text{sorted}}[-1]\}$
@@
Simultaneously, we enhance the image features within these critical regions. For each bounding box $B_{j}^{(i)}$ with a predicted class label $k$, the corresponding image features are amplified by a class-specific hyperparameter $\alpha_{k}$. This enhancement is applied to all pixels $p$ within the bounding box $B_{j}^{(i)}$ and across all $C$ feature channels. The operation is defined as:
@@
Here, the value of $\alpha_{k}$ is a class-specific hyperparameter, set based on the object’s typical size. The underlying principle is that smaller objects require a stronger feature boost to ensure their representation is not neglected during fusion. Consequently, classes like pedestrians and traffic cones are assigned a higher $\alpha_{k}$, whereas larger classes such as buses and trucks receive a more moderate value. This allows for a more targeted enhancement, tailoring the amplification intensity to the specific characteristics of each object category.
@@
Pixels where this discrepancy surpasses a predefined threshold $\tau$ are considered unreliable and subsequently masked out from $D_{\text{aligned}}^{(i)}$. This process yields a cleaner, but still sparse, depth map, $M^{(i)}$, where the value of a pixel at coordinates $(u,v)$ is given by:
@@
These two statistics are then broadcast to all pixels within their respective blocks, creating a densified depth map $D_{\text{dense}}^{(i)}$ and a densified gradient map $G_{\text{dense}}^{(i)}$. The visualization of these is shown in Figure 4, in which we also show the visualization of comparision between our method and others. Our method effectively addresses the misalignment between LiDAR and camera images. Furthermore, in regions with gradual depth changes, it successfully preserves relevant depth information without over-smoothing.
@@
The features generated by Discontinuity Aware Geometric Fusion (DAGF) guide final depth prediction, $\hat{D}^{(i)}$, from SGDM. The training is supervised by a composite loss function that leveraged both the dense map and the gradient map.
Method mAP NDS C T CV B Tr Ba M Bi P TC
BEVFusion-PKU (NeurIPS 22) (Liu et al. 2023) 67.9 71.0 87.3 59.8 28.9 73.5 41.2 73.8 74.6 59.8 85.4 68.2
BEVFusion-MIT (ICRA 23) (Liang et al. 2022) 68.5 71.4 88.2 61.7 30.2 75.1 41.5 72.5 76.3 64.2 87.5 81.0
Ours 71.5 73.6 89.8 68.5 35.1 77.2 45.5 78.0 80.5 68.3 90.1 82.0
@@
The per-pixel term is defined as $l_{\text{focal}}(u,v)=\text{FL}(\hat{D}^{(i)}(u,v),D_{\text{dense}}^{(i)}(u,v))$. Here, $\mathcal{V}$ is the set of all pixel coordinates with a valid (non-zero) depth, and $\text{FL}(\cdot,\cdot)$ denotes the Focal Loss function.
@@
As shown in Figure 2(b), our Structural Guidance Depth Modulator (SGDM) is designed for multi-modal depth estimation, intelligently integrating visual features from the camera with geometric data from our processed depth representation. The architecture first processes camera and depth features through parallel convolutional, batch normalization, and ReLU layers, while a key residual connection preserves the original camera feature stream. These encoded features are then concatenated and fed into a gated attention mechanism, where a dynamically generated spatial attention map, created via a convolutional path and a Sigmoid function modulating a parallel feature stream. Recognizing that the fusion process can dilute the rich semantic information inherent in the camera features, we introduce a residual connection. This architecture enables the network to recover and reinforce the original semantic context. The output of this module is a discrete probability distribution over a set of predefined depth bins for each pixel.
@@
Our experiments utilize large-scale multimodal autonomous driving dataset: nuScenes.The nuScenes dataset (Caesar et al. 2020) comprises 1,000 diverse urban scenes (approximately 20 seconds each) collected across Boston and Singapore. It features 1.4 million camera images, 390,000 LiDAR sweeps, 1.4 million RADAR sweeps, and 1.4 million annotated 3D bounding boxes across 23 object classes (e.g., vehicles, pedestrians, cyclists), supplemented with precise attributes (velocity, visibility) and HD maps. Evaluation follows the benchmark’s primary metrics: mean Average Precision (mAP) and the nuScenes Detection Score (NDS).
Implementation Details
Comparison Results
We evaluate our proposed framework against a comprehensive set of recent state-of-the-art (SOTA) 3D object detection methods on the challenging, large-scale benchmark: nuScenes. As detailed in Table 1, our method achieves a new state-of-the-art performance on the nuScenes validation set, reaching 71.5% mAP and 73.6% NDS with negligible increment of inference time. This result surpasses previous leading methods, including GraphBEV (Song et al. 2024a) (70.1% mAP, 72.9% NDS), BEVDiffuser (Ye et al. 2025) (69.2% mAP, 71.9% NDS), and the strong BEVFusion-PKU (Liu et al. 2023) baseline (67.9% mAP, 71.0% NDS). Our approach demonstrates consistently superior performance across the majority of object classes, notably achieving the highest AP scores for critical categories such as Car (89.8%), Truck (68.5%), Motorcycle (80.5%), and Pedestrian (90.1%). Significantly, we observe a substantial improvement in challenging classes like Construction Vehicle (35.1%) and Bicycle (68.3%), underscoring our model’s enhanced robustness in handling diverse and difficult-to-detect objects by effectively mitigating feature misalignment.
Ablation Study
@@
Table 2: Ablation study of our three proposed modules. The baseline uses none of our components. All other configurations use our SGDM module to process inputs from the enabled data-processing modules (PGDC, DAGF).
To further analyse component contributions, Table 3 presents a granular ablation. We incrementally test PGDC’s functions, the Depth Align Module (DAM) and the Feature Enhance Module (FEM), and DAGF’s representations ($D_{\text{dense}}$ and $G_{\text{dense}}$). The results show DAM provides the most significant initial boost. FEM offers a further gain, followed by the dense depth representation. Finally, the gradient representation pushes the model to its peak performance.
@@
Table 3: Granular ablation study. DAM: Depth Align Module. FEM: Feature Enhance Module. $D_{\text{dense}}$: Using the dense depth representation from DAGF. $G_{\text{dense}}$: Using the dense gradient representation from DAGF.
Finally, to isolate the impact of 2D detector accuracy, Table 4 shows an ablation study on 2D prior quality. We compare scenarios with no priors, noisy priors, realistic priors (YOLOv9), and perfect ground truth priors. The noisy condition refers to the YOLOv9 detection results with a synthetically introduced 10% false negative rate (missed detections) and a 10% false positive rate (incorrect detections).The results demonstrate that performance scales with the quality of the 2D priors, and our method with a realistic detector nearly closes the gap to the ceiling set by perfect information. For more ablation results about hyperparameters, please refer to the Appendix.
@@
Noisy 2D Priors 70.3 72.5
Realistic 2D Priors 71.5 73.6
@@
Table 4: Ablation study on impact of 2D detectors. We test scenarios with no 2D priors, noisy priors, realistic priors from YOLOv9, and perfect priors from ground truth.

### v2

@@ Abstract
Integrating LiDAR and camera inputs into a unified Bird’s-Eye-View (BEV) representation is crucial for enhancing 3D perception capabilities of autonomous vehicles. However, existing methods suffer from spatial misalignment between LiDAR and camera features, which causes inaccurate depth supervision in camera branch and erroneous fusion during cross-modal feature aggregation. The root cause of this misalignment lies in projection errors, stemming from calibration inaccuracies and rolling shutter effect. The key insight of this work is that locations of these projection errors are not random but highly predictable, as they are concentrated at object-background boundaries which 2D detectors can reliably identify. Based on this, our main motivation is to utilize 2D object priors to pre-align cross-modal features before fusion. To address local misalignment, we propose Prior Guided Depth Calibration (PGDC), which leverages 2D priors to alleviate misalignment and preserve correct cross-modal feature pairs. To resolve global misalignment, we introduce Discontinuity Aware Geometric Fusion (DAGF) to suppress residual noise from PGDC and explicitly enhance sharp depth transitions at object-background boundaries, yielding a structurally aware representation. To effectively utilize these aligned representations, we incorporate Structural Guidance Depth Modulator (SGDM), using a gated attention mechanism to efficiently fuse aligned depth and image features. Our method achieves SOTA performance on nuScenes validation dataset, with its mAP and NDS reaching 71.5% and 73.6% respectively.
Figure 1: Demonstration of projection errors under varying depth changes. LiDAR points from the distant wall (cyan) are incorrectly projected onto the foreground vehicle (which should appear red) due to a sharp depth change. In contrast, the boundary between background objects like the wall (light cyan) and garage (dark cyan) is correctly projected thanks to a gradual depth transition. This shows that misalignment is most severe at abrupt foreground-background boundaries.
@@ Introduction
Robust 3D perception is fundamental to autonomous driving, where effective sensor fusion is essential as different modalities offer complementary strengths. Images provide rich semantic information but lack accurate depth, while point clouds offer precise geometry and depth but are sparse and lack semantic context. Harnessing these complementary strengths while mitigating their limitations is key to building reliable perception systems.
To fully leverage complementary information from camera and LiDAR, current architectures either use LiDAR for explicit supervision of the 2D-to-3D transformation process (Li et al. 2023), or integrate two complementary information sources by fusing LiDAR and camera BEV representations (Liu et al. 2023; Liang et al. 2022). While effective, these methods are fundamentally challenged by the inherent spatial misalignment between sensors (Yu et al. 2023). Although the performance of these methods has a relatively high theoretical upper bound, reaching this bound requires perfect sensor alignment, which is frequently violated by projection errors. This discrepancy leads to two critical issues. First, it corrupts depth supervision signal, providing noisy or incorrect depth labels to image branch. Second, during cross-modal aggregation, spatial misalignment causes the fusion module to associate semantically mismatched image and geometric features, thereby degrading quality and reliability of the final fused representation.
@@
Existing approaches have attempted to mitigate cross-modal misalignment, yet each suffers from significant drawbacks. To ensure feature consistency, methods like transfusion (Bai et al. 2022) incorporate attention mechanism to query features of a specific modality, avoiding direct projection errors. While this effectively sidesteps the main cause of misalignment, it comes at the cost of sacrificing crucial contextual information. Other approaches like MetaBEV (Ge et al. 2023) and RobBEV (Wang et al. 2024) attempt to mitigate misalignment effects by designing more robust and adaptive fusion modules. Although these modules are more resilient to inconsistent inputs, they cannot correct the initial geometric errors from the 2D-3D view transformation. In essence, they are skillfully fusing features that have already been misplaced. Finally, global alignment techniques like GraphBEV (Song et al. 2024a) directly address the geometric problem, effectively eliminating misalignment in areas with steep depth gradients. However, they tend to unnecessarily smooth geometrically stable regions where misalignment is negligible or even entirely absent, thereby incorrectly modifying already-correct depth values.
Prevailing methods often overlook that misalignment is not randomly distributed, but originates from two primary sources: extrinsic calibration errors and motion-induced distortions. The resulting geometric projection errors are depth-dependent, being negligible for nearby objects but significantly exacerbated at greater depths. This depth-dependent distortion creates the most severe feature misalignment at boundaries between foreground objects and their backgrounds. Our approach is built upon the core insight that critical misalignments are concentrated at foreground-background boundaries where sharp discontinuities in depth occur (as illustrated in Figure 1), and we can leverage robust 2D object priors to identify these specific regions, locating real misalignment and accurately solving it.
@@
In this paper, we introduce three synergistic modules. First, to address the misalignment problem at its root, we propose (a) Prior Guided Depth Calibration (PGDC). Guided by the principle of “Look Before You Fuse”, our approach posits that rather than mitigating the effects of fused, misaligned data, it is more effective to proactively correct geometric inconsistencies using high-level semantic guidance before the fusion stage. PGDC actively uses 2D detection proposals as explicit geometric priors to locate and correct misaligned point cloud data, providing a significantly more accurate depth map. However, a corrected sparse map alone is not enough. To enable the network to fully understand the scene’s geometric structure, we introduce (b) Discontinuity Aware Geometric Fusion (DAGF). DAGF refines the PGDC-calibrated depth by masking points that exhibit large deviations from the raw depth map, subsequently replacing them with more reliable estimates, fitting the regions of sharp depth transition to the object-background boundaries. This self-correction mechanism corrects remaining misalignments when 2D priors is accurate, and reverts PGDC-induced over-smoothing when 2D priors are flawed. Eventually, DAGF produces a representation of the global depth structure. To effectively utilize these calibrated signals for view transformation, we introduce the (c) Structural Guidance Depth Modulator (SGDM). This module intelligently fuses image features and dense geometric representation from DAGF using a gated attention mechanism. Its purpose is to predict a highly accurate depth distribution for each pixel, enabling a more precise projection of features into the final BEV space.
@@
Figure 2: Overview of our proposed framework. (a) Prior Guided Depth Calibration (PGDC) and (b) Discontinuity Aware Geometric Fusion (DAGF), proactively mitigate multi-sensor feature misalignment before view transformation. And (c) Structural Guidance Depth Modulator (SGDM) intelligently fuses image features and dense geometric representation, predicting an accurate depth distribution. Finally, fusing rectified camera BEV features with LiDAR BEV features leads to robust 3D detection.
@@
Figure 3: The Prior Guided Depth Calibration (PGDC) module leverages 2D detection boxes as priors to precisely target and correct the most severe feature misalignments, which are caused by calibration errors and motion distortion. By applying localized smoothing to the point cloud within these detected regions, the module corrects the erroneous depth information. Simultaneously, it enhances the features of these critical image areas.
The most significant contributions stemming from our work are summarized below:
@@
  • We reveal that feature misalignment predominantly occurs at object-background boundaries and propose Prior Guided Depth Calibration (PGDC), a novel module that actively uses 2D priors to locate and resolve such misalignment, thereby providing more accurate depth information. To our best knowledge, our work is the first to use 2D detection results as explicit guidance for correcting cross-modal geometric inconsistencies before fusion.
@@
  • We introduce the Structural Guidance Depth Modulator (SGDM), an efficient attention-based module that fuses calibrated visual and geometric cues to generate an accurate depth distribution for view transformation.
@@
As shown in Figure 2, our proposed framework is a multi-sensor fusion pipeline built upon the strong BEVFusion (Liang et al. 2022) baseline, designed to explicitly mitigate feature misalignment in Bird’s Eye View (BEV) perception. The framework processes inputs from a set of $N$ surround-view cameras and a 360-degree LiDAR sensor. Our contribution is a combination of three synergistic modules. First, Prior Guided Depth Calibration (PGDC) uses 2D bounding boxes to correct misaligned LiDAR points at object boundaries, outputting a refined sparse depth map and enhanced image features. Second, Discontinuity Aware Geometric Fusion (DAGF) uses the corrected depth to generate a dense representation that captures reliable geometric structure. Finally, the Structural Guidance Depth Modulator (SGDM) fuses the enhanced image features (from PGDC) and the dense geometric representation (from DAGF) to predict a depth distribution for each camera view. Following the Lift-Splat-Shoot (Philion and Fidler 2020) paradigm, these are projected into a unified BEV feature map, which is then fused with LiDAR BEV features to produce robust 3D detection results.
@@
As shown in Figure 3, the Prior Guided Depth Calibration (PGDC) module operates independently on each of the $N$ camera views to correct the initial sparse depth supervision derived from LiDAR. For each view $i\in\{1,...,N\}$, the inputs are the image features $F_{\text{img}}^{(i)}\in\mathbb{R}^{H\times W\times C}$ and the corresponding sparse depth map $D_{\text{raw}}^{(i)}\in\mathbb{R}^{H\times W}$, which is generated by projecting the LiDAR point cloud into that camera’s image plane. To better explain this module, we have divided it into the 2D Guided Depth Align Module and the 2D Camera Features Enhance Module.
First, a 2D detection head provides a set of bounding boxes $\{B_{j}^{(i)}\}$ for each image. In 2D Guided Depth Align Module, based on our observation that misalignment is concentrated at object boundaries, we use these boxes to isolate critical regions. For each box $B_{j}^{(i)}$, we filter the LiDAR points whose projections fall within it. To mitigate misalignment, we introduce a novel smoothing operation detailed in Algorithm 1. Instead of simple averaging, our method captures the local depth structure more effectively. For each LiDAR point projected to a pixel $p$ with depth $d_{p}$, we first use a KD-Tree to find its 10 nearest neighbors, $\mathcal{N}_{p}$. We then select the two neighbors with the smallest depth and the two with the largest depth, forming a set of four critical neighbors, $\mathcal{N}_{\text{critical}}\subset\mathcal{N}_{p}$. This selection strategy is designed to simultaneously capture the object’s own depth consistency (via the nearest points) and the sharp depth discontinuity at the boundary between the object and the background (via the farthest points), thereby preserving critical information while smoothing noise. The depth of the original point is concatenated with the depths of these four selected neighbors, creating a 5-channel feature map, $f_{p}$, for that pixel:
@@
This feature map is then processed through a lightweight convolutional block to produce the final single-channel, smoothed depth value:
@@
9:  $\mathcal{D}_{\text{critical}}\leftarrow\mathcal{D}_{\text{sorted}}[0:2]\cup\mathcal{D}_{\text{sorted}}[-2:]$
@@
Simultaneously, in 2D Camera Features Enhance Module, we enhance the image features within these critical regions. For each bounding box $B_{j}^{(i)}$ with a predicted class label $k$, the corresponding image features are amplified by a class-specific hyperparameter $\alpha_{k}$. This enhancement is applied to all pixels $p$ within the bounding box $B_{j}^{(i)}$ and across all $C$ feature channels. The operation is defined as:
@@
where the value of $\alpha_{k}$ is a class-specific hyperparameter, set based on the object’s typical size. The underlying principle is that smaller objects require a stronger feature boost to ensure their representation is not neglected during fusion. Consequently, small classes like pedestrians and traffic cones are assigned a higher $\alpha_{k}$, whereas larger classes such as buses and trucks receive a more moderate value. This allows for a more targeted enhancement, tailoring the amplification intensity to the specific characteristics of each object category.
@@
Method mAP NDS C T CV B Tr Ba M Bi P TC
BEVFusion-PKU (NeurIPS 22) (Liang et al. 2022) 67.9 71.0 87.3 59.8 28.9 73.5 41.2 73.8 74.6 59.8 85.4 68.2
BEVFusion-MIT (ICRA 23) (Liu et al. 2023) 68.5 71.4 88.2 61.7 30.2 75.1 41.5 72.5 76.3 64.2 87.5 81.0
Ours 71.5 73.6 89.8 68.5 35.1 77.2 45.5 78.0 80.5 68.3 90.1 82.0
@@
Pixels where this discrepancy surpasses a threshold $\tau$, defined as 10% of the pixel’s value in $D_{\text{raw}}^{(i)}$, are considered unreliable and subsequently masked out. This process yields a cleaner, but still sparse, depth map, $M^{(i)}$, where the value of a pixel at coordinates $(u,v)$ is given by:
@@
These two statistics are then broadcast to all pixels within their respective blocks, creating a densified depth map $D_{\text{dense}}^{(i)}$ and a densified gradient map $G_{\text{dense}}^{(i)}$. The visualization of these maps is shown in Figure 4, in which we also show the visualization of comparison between our method and others. Our method effectively addresses the misalignment between LiDAR and camera images. Furthermore, in regions with gradual depth changes, it successfully preserves correct depth information while GraphBEV (Song et al. 2024a) falsely modifies already-correct depth values.
@@
The features generated by Discontinuity Aware Geometric Fusion (DAGF) guide final depth prediction, $\hat{D}^{(i)}$, from SGDM. The training is supervised by a composite loss function that leverages both the dense map and the gradient map.
@@
The per-pixel term is defined as $l_{\text{focal}}(u,v)=\text{FL}(\hat{D}^{(i)}(u,v),D_{\text{dense}}^{(i)}(u,v))$. Here, $\mathcal{V}$ is the set of all pixel coordinates with a valid (non-zero) depth, and $\text{FL}(\cdot,\cdot)$ denotes the Focal Loss function. The hyperparameters for the Focal Loss function are set to $\gamma=2.0$ and $\alpha=0.25$.
@@
As shown in Figure 2(c), our Structural Guidance Depth Modulator (SGDM) is designed for multi-modal depth estimation, intelligently integrating visual features from the camera with geometric data from our processed depth representation. The architecture first processes camera and depth features through parallel convolutional layers in order to extract and normalize modality-specific features. These encoded features are then concatenated and fed into a processing block where a gated attention mechanism generates a spatial attention map. This map modulates the initial depth prediction, effectively learning a confidence score for each pixel’s placement in 3D space.
Recognizing that the fusion process can dilute the rich semantic information inherent in the camera features, we introduce a crucial residual connection to preserve the original camera feature stream. This connection acts as a direct information pathway, carrying the pristine visual features forward and bypassing the fusion block.
This architecture thus creates a powerful synergy: the residual path guarantees the preservation of critical visual context, while the attention gate intelligently modulates the feature map, allowing the network to selectively emphasize reliable information. The final output of this module is a discrete probability distribution over a set of predefined depth bins for each pixel, framing depth estimation as a more stable, per-pixel classification task.
@@
Our experiments use the large-scale multimodal autonomous driving dataset, nuScenes (Caesar et al. 2020), which contains 1,000 diverse urban scenes ( 20 seconds each) collected in Boston and Singapore. It includes 1.4M camera images, 390K LiDAR sweeps, 1.4M RADAR sweeps, and 1.4M annotated 3D bounding boxes spanning 23 object classes (e.g., vehicles, pedestrians, cyclists), along with rich attributes (e.g., velocity, visibility) and HD maps. We evaluate using the benchmark’s main metrics: mean Average Precision (mAP) and nuScenes Detection Score (NDS).
@@
Table 2: Ablation study of our three proposed modules. The baseline uses none of our components.
Implementation Details
Comparison Results
We evaluate our proposed framework against a comprehensive set of recent state-of-the-art (SOTA) 3D object detection methods on the challenging, large-scale benchmark: nuScenes. As detailed in Table 1, our method achieves a new state-of-the-art performance on the nuScenes validation set, reaching 71.5% mAP and 73.6% NDS with negligible increment of inference time. This result surpasses previous leading methods, including GraphBEV (Song et al. 2024a) (70.1% mAP, 72.9% NDS), BEVDiffuser (Ye et al. 2025) (69.2% mAP, 71.9% NDS), and the strong BEVFusion-PKU (Liu et al. 2023) baseline (67.9% mAP, 71.0% NDS).
Ablation Study
To further analyze component contributions, Table 3 presents a granular ablation. We incrementally test PGDC’s functions, the Depth Align Module (DAM) and the Feature Enhance Module (FEM), and DAGF’s representations ($D_{\text{dense}}$ and $G_{\text{dense}}$). The results show DAM provides the most significant initial boost. FEM offers a further gain, followed by the dense depth representation. Finally, the gradient representation pushes the model to its peak performance.
@@
Table 3: Granular ablation study. We seperate PGDC into DAM: Depth Align Module and FEM: Feature Enhance Module.
@@
Full-Image Prior 69.4 71.8
YOLO-X Priors 70.3 72.5
Realistic 2D Priors (YOLOv9) 71.5 73.6
@@
Table 4: Ablation study on the impact of 2D detectors.
Finally, to isolate the impact of 2D detector accuracy, Table 4 shows an ablation study on 2D prior quality. The results demonstrate that while performance generally scales with the quality of the 2D priors, our method is notably robust. Even with completely random priors, the model’s performance is not significantly harmed. Furthermore, using a coarse prior that covers the full image still brings an improvement over using no prior at all. The system maintains structural accuracy even when faulty 2D priors cause the PGDC to falsely over-smooth the correct depth. This is because after the DAGF module removes the incorrectly smoothed points, refilling the boundary regions accurately preserves the point cloud’s structure.
A more comprehensive ablation study, including hyperparameter sensitivity and the influence of 2D detectors on our proposed modules, can be found in the Appendix.